OpenStudy (anonymous):

what rule is -1-2=1?

OpenStudy (katrinakaif):

That is not quite possible. -1 - 2 = -3 not +1

OpenStudy (anonymous):

not quite...... impossible*

OpenStudy (anonymous):

\[(-1)-2\neq1\]

OpenStudy (anonymous):

"that is not quite possible" inferring that is possible? lolol

OpenStudy (anonymous):

maybe you mean to say -1-(-2)=1

OpenStudy (anonymous):

but my teacher said it was possible

OpenStudy (anonymous):

did you miss a negative sign? when posting the question?

OpenStudy (anonymous):

no

OpenStudy (katrinakaif):

You may have written the problem incorrectly.

OpenStudy (anonymous):

oh, well then your teacher lied!

OpenStudy (anonymous):

because believe me when i say -1-2=-3 not 1

OpenStudy (paxpolaris):

\[-1-2 = -3\]\[-1--2=1\]

Directrix (directrix):

@Shiny --> Would you check and see it the problem is written correctly?

OpenStudy (anonymous):

ITS WRITTEN CORRECTLY

Directrix (directrix):

Okay, let's look at it: what rule is -1-2=1? The equation reads in words as negative one minus two equals one. That is not true because negative one minus two equals negative one. Because the statement is false is why there is no rule.

Directrix (directrix):

And, that's why I asked you to check the problem. It is so easy to type the wrong number or symbol.